# Space spanned by characteristic exponents

### v2

@@ 3. An equidistribution theorem
Results of equidistribution-type play an important role in the study of arithmetic geometry and arithmetic dynamics in past decades. In this section, we state and prove a reformulation (Theorem 3.1) of Yuan’s equidistribution theorem [32, Theorem 3.1]. The reason that we present the theorem is that the equidistribution theorems in the literature (for example, [32]) are only dealing with the Galois orbit of a single small point at one time, which is not sufficient for our proof of Theorem 1.5. Our statement is slightly stronger than [32, Theorem 3.1] as we work with a sequence of (weighted) countable Galois invariant subsets, and this is useful for applications in dynamics. The presented proof of Theorem 3.1 is by applying Yuan’s original equidistribution theorem [32, Theorem 3.1]. We follow the terminology in [32].
For the convenience of readers, we present a short review of some basic notions in Arakelov geometry. See [36, 32, 20, 33] for more details.
Let $K$ be a number field, $X$ a projective variety of dimension $n-1$ over $K$, and $\mathcal{L}$ a line bundle on $X$. We fix an embedding $\overline{K}\to\mathbb{C}_{v}$ for each place $v\in\mathcal{M}_{K}$ of $K$. An adelic metric on $\mathcal{L}$ is a $\mathbb{C}_{v}$-norm $\|\cdot\|_{v}$ on the fiber $\mathcal{L}_{\mathbb{C}_{v}}(x)$ for every $x\in X(\overline{K})$ and $v\in\mathcal{M}_{K}$, satisfying certain continuity and coherence conditions, see [36, §1] for a precise definition.
An adelic metric on $\mathcal{L}$ is called an algebraic metric (or a model metric) if it is induced by an $\mathcal{O}_{K}$-model $(\mathcal{X},\tilde{\mathcal{L}})$ of $(X,\mathcal{L}^{e})$ for some $e\in\mathbb{Z}_{>0}$ (i.e., $\mathcal{X}$ is an integral scheme projective and flat over $\mathcal{O}_{K}$ and $\tilde{\mathcal{L}}$ is a Hermitian line bundle over $\mathcal{X}$ such that the generic fiber of $(\mathcal{X},\tilde{\mathcal{L}})$ is $(X,\mathcal{L}^{e})$). Here the induced metric at an archimedean place is induced by the Hermitian metric of $\tilde{\mathcal{L}}$, and for a non-archimedean place $v$ and $x\in X(\overline{K})$, the induced metric on $\mathcal{L}_{\mathbb{C}_{v}}(x)$ is determined by the lattice $(\tilde{x}^{*}\tilde{\mathcal{L}}_{\mathcal{O}_{\mathbb{C}_{v}}})^{1/e}$ in $\mathcal{L}_{\mathbb{C}_{v}}(x)$, where $\tilde{x}:{\rm Spec}(\mathcal{O}_{\mathbb{C}_{v}})\to\mathcal{X}_{\mathcal{O}_{\mathbb{C}_{v}}}$ is the extension of $x$. The metric induced by such a model always satisfies the conditions of an adelic metric, see [36, §1]. The algebraic metric is called semipositive if $\tilde{\mathcal{L}}$ has semipositive curvature at all archimedean places and non-negative degree on every vertical curve of $\mathcal{X}$.
An (adelic) metric on $\mathcal{L}$ is called semipositive if it is the uniform limit of some sequence of semipositive algebraic metrics on $\mathcal{L}$. Assume that $\mathcal{L}$ is ample and equip it with a semipositive (adelic) metric (such a metric always exists). The resulting metrized line bundle is denoted by $\overline{\mathcal{L}}$. Then for every closed subvariety $Y$ of $X_{\overline{K}}$, we can define its $\overline{\mathcal{L}}$-height $h_{\overline{\mathcal{L}}}(Y)$ by
@@
where $\overline{Y}$ is the image of $Y\to X_{\overline{K}}\to X$ (i.e., the closed subvariety of $X$ corresponding to the Galois orbit of $Y$), and $\hat{c}_{1}(\overline{\mathcal{L}}|_{\overline{Y}})^{\dim(Y)+1}$ is the arithmetic self-intersection number of ¯$\overline{\mathcal{L}}|_{\overline{Y}}$ on $\overline{Y}$. In particular, for every algebraic point $x\in X(\overline{K})$, we get its height $h_{\overline{\mathcal{L}}}(x)$ and the function $h_{\overline{\mathcal{L}}}:{X(\overline{K})}\to\mathbb{R}$ is a Weil height associated to $\mathcal{L}$ (up to a normalizing factor $[K:\mathbb{Q}]$). For an algebraic point $x\in X(\overline{K})$, we have
@@
Let $K\subset\overline{\mathbb{Q}}$ be a number field and $X$ be a projective variety over $K$. Fix an embedding of $\overline{\mathbb{Q}}$ into $\mathbb{C}$. Let $\overline{\mathcal{L}}$ be a metrized line bundle on $X$ such that $\mathcal{L}$ is ample and the metric is semipositive. Let $\mu:=\deg_{\mathcal{L}}(X)^{-1}c_{1}(\overline{\mathcal{L}})^{\dim(X)}_{\mathbb{C}}$ be the canonical probability measure on $X(\mathbb{C})$ associated to $\overline{\mathcal{L}}$. For $n\in\mathbb{Z}_{>0}$, let $S_{n}$ be a countable subset of $X(\overline{\mathbb{Q}})$ which is ${\rm Gal}(\overline{\mathbb{Q}}/K)$-invariant. For every $n\geq 1$ and $y\in S_{n}$, fix a real number $a_{n,y}>0$ such that $\sum_{y\in S_{n}}a_{n,y}=1$ and $a_{n,y}=a_{n,\sigma(y)}$, for all $n\geq 1$, $y\in S_{n}$, and $\sigma\in{\rm Gal}(\overline{\mathbb{Q}}/K)$. Assume that the following two conditions hold:
@@
(2) (generic) for every proper subvariety $V\subsetneqq X$ of $X$, $\sum_{y\in S_{n}\cap V}a_{n,y}\to 0$ as $n\to+\infty$.
Then the measure $\mu_{n}:=\sum_{y\in S_{n}}a_{n,y}\delta_{y}$ converges weakly to $\mu$ on $X(\mathbb{C})$ as $n\to+\infty$, where $\delta_{y}$ denotes the Dirac measure at the point $y$, i.e., for all continuous function $g$ on $X(\mathbb{C})$ (with the complex topology), we have
[…]
@@
In [32, Theorem 3.1], Yuan prove the theorem for the case where every $S_{n}$ ($n\geq 1$) is consisting of a single Galois orbit, i.e., $m_{n}=1$ for every $n\geq 1$.
[…]
@@
As $a_{n,y}$ is constant for $y$ in its Galois orbit, we can define $a_{n,i}:=a_{n,y}$ where $y$ is an arbitrary point in $O_{n,i}$, for all $n\geq 1$ and $1\leq i\leq m_{n}$. For all $n\geq 1$ and $1\leq i\leq m_{n}$, set $\mu_{n,i}:=\sum_{y\in O_{n,i}}\delta_{y}$. Then we have
$\sum_{i=1}^{m_{n}}a_{n,i}(\#O_{n,i})=1\text{ and }\mu_{n}=\sum_{i=1}^{m_{n}}a_{n,i}\mu_{n,i}$
for every $n\geq 1$. As $h_{\overline{\mathcal{L}}}(y)$ is constant for $y$ in its Galois orbit, we can define $h_{n,i}:=h_{\overline{\mathcal{L}}}(y)$ where $y$ is an arbitrary point in $O_{n,i}$, for all $n\geq 1$ and $1\leq i\leq m_{n}$.
The small condition $(1)$ asserts that $\sum_{i=1}^{m_{n}}a_{n,i}(\#O_{n,i})h_{n,i}\to h_{\overline{\mathcal{L}}}(X)$ as $n\to+\infty$. Taking a suitable subsequence, we may assume that
@@
As $X$ has only countably many subvarieties defined over $\overline{\mathbb{Q}}$, there is a sequence $V_{1}\subseteq V_{2}\subseteq\cdots\subseteq V_{n}\subseteq\cdots\subsetneqq X$ of increasing proper subvarieties of $X$ defined over $K$, such that for every proper subvariety $V$ of $X$ defined over $\overline{\mathbb{Q}}$, there is $N=N(V)\in\mathbb{Z}_{>0}$ such that $V\subseteq V_{n}$ for all integers $n\geq N$. (By considering the Galois orbits of subvarieties, we can assume that the $V_{n}$’s are defined over $K$.) By Zhang’s theorem of successive minimum [36, Theorem 1.10], we may further assume that
@@
The generic condition $(2)$ shows that after replacing $(S_{n})$ by a suitable subsequence, we may assume that
@@
$\frac{1+M}{n^{2}}$ $\geq\sum_{i\in G_{n}}a_{n,i}(\#O_{n,i})(h_{n,i}-h_{\overline{\mathcal{L}}}(X))\geq\frac{-\sum_{i\in H_{n}}a_{n,i}(\#O_{n,i})}{n^{2}}+\frac{\sum_{i\in I_{n}}a_{n,i}(\#O_{n,i})}{n}$
@@
[…]
@@
for every $n\geq 1$, i.e., the small condition $(1)$ holds for $(O_{n,i_{n}})_{n}$. Applying Yuan’s equidistribution Theorem [32, Theorem 3.1] to $(O_{n,i_{n}})_{n}$, we have $\mu_{n,i_{n}}\to\mu$ as $n\to+\infty$. Thus we have
[…]
@@
For a place $v\in\mathcal{M}_{K}$, we can consider two canonical analytic $v$-spaces of $X$, i.e., the $\mathbb{C}_{v}$-analytic space $X^{an}_{\mathbb{C}_{v}}$ and the $K_{v}$-analytic space $X^{an}_{K_{v}}$. We call the former the geometric case and the latter the algebraic case, see [32, §3.1]. Theorem 3.1 deals with the geometric case at an archimedean place. The same idea also applies for a non-archimedean place or the algebraic case, which gives the full analogy of [32, Theorems 3.1 and 3.2].
In order to check the “generic” condition in Theorem 3.1, we need the following lemma. The proof uses ergodic theory with respect to the constructible topology (on algebraic varieties) introduced by Xie [28].
@@
Let $K$ be a number field and $X$ be a projective variety over $K$. Given a dominant endomorphism $f:X\to X$, and a sequence $(x_{n})_{n=1}^{\infty}$ of $f$-periodic points in $X(\overline{K})$ which are pairwise distinct. Assume that $(x_{n})_{n=1}^{\infty}$ is generic in $X$, i.e., every proper Zariski closed subset $Z\subsetneqq X$ contains only finitely many $x_{n}$. Then for every proper subvariety $V\subsetneqq X$, we have
(3.11) #$\frac{\#(V\cap O_{f}(x_{n}))}{\#O_{f}(x_{n})}\to 0,\ \text{as}\ n\to+\infty,$
@@
#$\frac{\#(V\cap O_{f}(x_{n_{k_{l}}}))}{\#O_{f}(x_{n_{k_{l}}})}\to 0,\ \text{as}\ l\to+\infty.$
Fix a proper subvariety $V$ in $X$. Let $\lvert X\rvert$ be $X$ equipped with the constructible topology (i.e., the topology of $X$ generated by all its Zariski closed and open subsets) and $\mathcal{M}^{1}(\lvert X\rvert)$ be the space of all probability Radon measures on $\lvert X\rvert$ with the topology of weak convergence relative to all continuous functions on $\lvert X\rvert$. Then $\mathcal{M}^{1}(\lvert X\rvert)$ is sequentially compact (cf. [28, Corollary 1.14]). For $n\in\mathbb{Z}_{>0}$, set
@@
Denote the characteristic function of $V\subsetneqq\lvert X\rvert$ by $1_{V}$, then $1_{V}$ is continuous with respect to the constructible topology. By the sequential compactness of $\mathcal{M}^{1}(\lvert X\rvert)$, it suffices to show that for every subsequence $(n_{k})_{k}$ of $(n)_{n=1}^{\infty}$ with $m_{n_{k}}\to m\ \text{as}\ k\to+\infty$ in $\mathcal{M}^{1}(\lvert X\rvert)$ for some $m\in\mathcal{M}^{1}(\lvert X\rvert)$, we have
@@
Without loss of generality, we may assume that $(m_{n})$ itself converges to a measure $m\in\mathcal{M}^{1}(\lvert X\rvert)$; and it suffices to show (3.12) in this case. As $f_{*}m_{n}=m_{n}$ for all $n\geq 1$, we see that $f_{*}m=m$ as well. Then according to [28, Lemma 5.3], $m$ must have the form $m=\sum_{y\in S}a_{y}\delta_{O_{f}(y)}$, where $S$ is a countable set of $f$-periodic points in $\lvert X\rvert$ (note that the points in $S$ may not be closed with respect to the Zariski topology of $X$), $a_{y}\in\mathbb{R}_{>0}$ with $\sum_{y\in S}a_{y}=1$, and $\delta_{O_{f}(y)}=(\#O_{f}(y))^{-1}\sum_{z\in O_{f}(y)}\delta_{z}$ for $y\in S$. As $m_{n}\to m$, we get
@@
Suppose that (3.12) fails. Then there must be a scheme point $y\in S$ with $a_{y}>0$ and $V\cap O_{f}(y)\neq\emptyset$. Denote the exact period of $y$ under $f$ by $k$. Let $Y$ be the Zariski closure of $\{y\}$. Then $Y\subseteq\cup_{j=0}^{k-1}f^{j}(V)$, hence $Y$ is also a proper Zariski closed subset of $X$. Note that
[…]
Hence for every sufficiently large integer $n\gg 1$, we have
@@
where the last equality follows from $f^{k}(y)=y$ and $Y=\overline{\{y\}}^{{\rm Zar}}$. Note that
@@
hence $\cup_{j=0}^{k-1}f^{j}(Y)$ contains only finitely many $x_{n}$ because $(x_{n})$ is generic, contradicting to (3.13). ∎
@@
Recall that $f:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ is a rational map of degree $d\geq 2$ and there is a number field $K$ such that for every $z_{0}\in{\rm Per}(f)$, we have $\lvert\rho_{f}(z_{0})\rvert^{n}\in K$ for some integer $n=n(z_{0})\geq 1$. We need to show that such a $f$ must be exceptional (Definition 1.1).
Assume by contradiction that $f$ is not exceptional. In particularly, $f$ is not a flexible Lattès map. For every $z_{0}\in{\rm Per}(f)$, we have $\lvert\rho_{f}(z_{0})\rvert\in\overline{\mathbb{Q}}$, thus $f$ is defined over $\overline{\mathbb{Q}}$ (hence over some number field) by Theorem 1.2. After replacing $K$ by a suitable finite extension of $K$, we may assume that both $f$ and $\overline{f}$ are defined over $K$. Here we denote by $\overline{f}$ the rational map obtained from $f$ via replacing the coefficients by their complex conjugations. According to [10, Theorem 9 and Lemma 11], which are essentially proved by Zdunik [35], there exists a sequence $(x_{n})_{n=1}^{\infty}$ of distinct points in ${\rm Per}^{*}(f)$ such that
@@
Clearly, $\mathcal{L}_{f}=\mathcal{L}_{\overline{f}}$. For an arbitrary $x\in{\rm Per}(f)$, we have
$\overline{x}\in{\rm Per}(\overline{f}),n_{f}(x)=n_{\overline{f}}(\overline{x}),\text{ and }\rho_{f}(x)=\overline{\rho_{\overline{f}}(\overline{x})},$
hence $\chi_{f}(x)=\chi_{\overline{f}}(\overline{x})$.

### v3

@@ 3. An equidistribution theorem
Equidistribution-type results have played a significant role in arithmetic geometry and arithmetic dynamics over the past decades. In this section, we state and prove a reformulation (Theorem 3.1) of Yuan’s equidistribution theorem [32, Theorem 3.1]. We present this theorem because existing equidistribution theorems in the literature (e.g., [32]) typically handle the Galois orbit of a single small point at a time, which is insufficient for our proof of Theorem 1.5. Our statement is slightly stronger than [32, Theorem 3.1] as we work with sequences of (weighted) countable Galois-invariant subsets, which is particularly useful for applications in dynamics. The proof of Theorem 3.1 is based on applying Yuan’s original equidistribution theorem [32, Theorem 3.1]. We follow the terminology in [32].
For the convenience of readers, we briefly review some basic notions in Arakelov geometry. More details can be found in [36, 32, 20, 33].
Let $K$ be a number field, $X$ a projective variety of dimension n1n-1 over $K$, and $\mathcal{L}$ a line bundle on $X$. We fix an embedding $\overline{K}\to\mathbb{C}_{v}$ for each place $v\in\mathcal{M}_{K}$ of $K$. An adelic metric on $\mathcal{L}$ is a $\mathbb{C}_{v}$-norm $\|\cdot\|_{v}$ on the fiber $\mathcal{L}_{\mathbb{C}_{v}}(x)$ for every $x\in X(\overline{K})$ and $v\in\mathcal{M}_{K}$, satisfying certain continuity and coherence conditions; see [36, §1] for a precise definition.
[…]
An (adelic) metric on $\mathcal{L}$ is called semipositive if it is the uniform limit of a sequence of semipositive algebraic metrics on $\mathcal{L}$. Assume that $\mathcal{L}$ is ample and equipped with a semipositive (adelic) metric (which always exists). The resulting metrized line bundle is denoted by $\overline{\mathcal{L}}$. Then for every closed subvariety $Y$ of $X_{\overline{K}}$, we can define its $\overline{\mathcal{L}}$-height $h_{\overline{\mathcal{L}}}(Y)$ by
@@
where $\overline{Y}$ is the image of $Y\to X_{\overline{K}}\to X$ (i.e., the closed subvariety of $X$ corresponding to the Galois orbit of $Y$), and $\hat{c}_{1}(\overline{\mathcal{L}}|_{\overline{Y}})^{\dim(Y)+1}$ is the arithmetic self-intersection number of $\overline{\mathcal{L}}|_{\overline{Y}}$ on $\overline{Y}$. In particular, for every algebraic point $x\in X(\overline{K})$, we get its height $h_{\overline{\mathcal{L}}}(x)$, and the function $h_{\overline{\mathcal{L}}}:{X(\overline{K})}\to\mathbb{R}$ is a Weil height associated to $\mathcal{L}$ (up to a normalizing factor $K:\mathbb{Q}]$). For an algebraic point $x\in X(\overline{K})$, we have
@@
Let $K\subset\overline{\mathbb{Q}}$ be a number field and $X$ be a projective variety over $K$. Fix an embedding of $\overline{\mathbb{Q}}$ into $\mathbb{C}$. Let $\overline{\mathcal{L}}$ be a metrized line bundle on $X$ such that $\mathcal{L}$ is ample and the metric is semipositive. Let $\mu:=\deg_{\mathcal{L}}(X)^{-1}c_{1}(\overline{\mathcal{L}})^{\dim(X)}_{\mathbb{C}}$ be the canonical probability measure on $X(\mathbb{C})$ associated to $\overline{\mathcal{L}}$. For $n\in\mathbb{Z}_{>0}$, let $S_{n}$ be a countable subset of $X(\overline{\mathbb{Q}})$ which is ${\rm Gal}(\overline{\mathbb{Q}}/K)$-invariant. For every $n\geq 1$ and $y\in S_{n}$, fix a real number $a_{n,y}>0$ such that $\sum_{y\in S_{n}}a_{n,y}=1$ and $a_{n,y}=a_{n,\sigma(y)}$ for all $n\geq 1$, $y\in S_{n}$, and $\sigma\in{\rm Gal}(\overline{\mathbb{Q}}/K)$. Assume that the following two conditions hold:
@@
(2) (generic) for every proper subvariety $V\subsetneqq X$, $\sum_{y\in S_{n}\cap V}a_{n,y}\to 0$ as $n\to+\infty$.
Then the measure $\mu_{n}:=\sum_{y\in S_{n}}a_{n,y}\delta_{y}$ converges weakly to $\mu$ on $X(\mathbb{C})$ as $n\to+\infty$, where $\delta_{y}$ denotes the Dirac measure at the point $y$; i.e., for every continuous function $g$ on $X(\mathbb{C})$ (with the complex topology), we have
(3.5) $\lim\limits_{n\to\infty}\sum_{y\in S_{n}}a_{n,y}g(y)=\int_{X(\mathbb{C})}gd\mu.$
@@
In [32, Theorem 3.1], Yuan proved the theorem for the case where each $S_{n}$ ($n\geq 1$) consists of a single Galois orbit, i.e., $m_{n}=1$ for every $n\geq 1$.
By contradiction, after passing to a subsequence of $(S_{n})$, we may assume that there is a continuous (real-valued) function $g$ on $X(\mathbb{C})$ and a constant $c\in\mathbb{R}_{>0}$ such that $\int_{X(\mathbb{C})}gd\mu=0$ and
(3.6) $\int_{X(\mathbb{C})}gd\mu_{n}\geq c>0$
@@
As $a_{n,y}$ is constant on Galois orbits, we can define $a_{n,i}:=a_{n,y}$ for an arbitrary $y\in O_{n,i}$, for all $n\geq 1$ and $1\leq i\leq m_{n}$. For all $n\geq 1$ and $1\leq i\leq m_{n}$, set $\mu_{n,i}:=\sum_{y\in O_{n,i}}\delta_{y}$. Then we have
$\sum_{i=1}^{m_{n}}a_{n,i}(\#O_{n,i})=1\quad\text{and}\quad\mu_{n}=\sum_{i=1}^{m_{n}}a_{n,i}\mu_{n,i}$
for every $n\geq 1$. As $h_{\overline{\mathcal{L}}}(y)$ is constant on Galois orbits, we can define $h_{n,i}:=h_{\overline{\mathcal{L}}}(y)$ for an arbitrary $y\in O_{n,i}$, for all $n\geq 1$ and $1\leq i\leq m_{n}$.
The small condition (1) implies that $\sum_{i=1}^{m_{n}}a_{n,i}(\#O_{n,i})h_{n,i}\to h_{\overline{\mathcal{L}}}(X)$ as $n\to+\infty$. After taking a suitable subsequence, we may assume that
@@
As $X$ has only countably many subvarieties defined over $\overline{\mathbb{Q}}$, there is an increasing sequence $V_{1}\subseteq V_{2}\subseteq\cdots\subseteq V_{n}\subseteq\cdots\subsetneqq X$ of proper subvarieties of $X$ defined over $K$ such that for every proper subvariety $V$ of $X$ defined over $\overline{\mathbb{Q}}$, there exists $N=N(V)\in\mathbb{Z}_{>0}$ with $V\subseteq V_{n}$ for all integers $n\geq N$. (By considering the Galois orbits of subvarieties, we may assume that the $V_{n}$ are defined over $K$.) By Zhang’s theorem of successive minima [36, Theorem 1.10], we may further assume that
@@
The generic condition (2) implies that after passing to a subsequence of $(S_{n})$, we may assume that
@@
$\frac{1+M}{n^{2}}$ $\geq\sum_{i\in G_{n}}a_{n,i}(\#O_{n,i})(h_{n,i}-h_{\overline{\mathcal{L}}}(X))\geq-\frac{\sum_{i\in H_{n}}a_{n,i}(\#O_{n,i})}{n^{2}}+\frac{\sum_{i\in I_{n}}a_{n,i}(\#O_{n,i})}{n}$
@@
Set $\mu_{n}^{\prime}:=\sum_{i\in H_{n}}a_{n,i}\mu_{n,i}$ for $n\geq 1$. Then $\mu_{n}-\mu_{n}^{\prime}$ is a (positive) measure on $X(\mathbb{C})$ whose total mass tends to 0 as $n\to+\infty$. After removing finitely many terms, we may assume that $\int_{X(\mathbb{C})}gd\mu_{n}^{\prime}\geq c/2$ for every $n\geq 1$, since $g$ is bounded on the compact set $X(\mathbb{C})$. Hence there is $i_{n}\in H_{n}$ such that $\int_{X(\mathbb{C})}gd\mu_{n,i_{n}}\geq c/2$; fix such an $i_{n}$ for each $n\geq 1$. Since $O_{n,i_{n}}\cap V_{n}=\emptyset$, the sequence $(O_{n,i_{n}})_{n}$ of Galois orbits satisfies the generic condition (2) by construction of $(V_{n})$. By (3.8) and the definition of $H_{n}$, we have
@@
for every $n\geq 1$, i.e., the small condition (1) holds for $(O_{n,i_{n}})_{n}$. Applying Yuan’s equidistribution theorem [32, Theorem 3.1] to $(O_{n,i_{n}})_{n}$, we get $\mu_{n,i_{n}}\to\mu$ as $n\to+\infty$. Thus,
$0<\frac{c}{2}\leq\lim_{n\to\infty}\int_{X(\mathbb{C})}gd\mu_{n,i_{n}}=\int_{X(\mathbb{C})}gd\mu=0,$
@@
For a place $v\in\mathcal{M}_{K}$, we can consider two canonical analytic $v$-spaces of $X$: the $\mathbb{C}_{v}$-analytic space $X^{an}_{\mathbb{C}_{v}}$ and the $K_{v}$-analytic space $X^{an}_{K_{v}}$. We refer to the former as the geometric case and the latter as the algebraic case; see [32, §3.1]. Theorem 3.1 deals with the geometric case at an archimedean place. The same idea applies to non-archimedean places or the algebraic case, which gives the full analogy of [32, Theorems 3.1 and 3.2].
To verify the “generic” condition in Theorem 3.1, we need the following lemma. The proof uses ergodic theory with respect to the constructible topology (on algebraic varieties) introduced by Xie [28].
@@
Let $K$ be a number field and $X$ be a projective variety over $K$. Given a dominant endomorphism $f:X\to X$ and a sequence $(x_{n})_{n=1}^{\infty}$ of pairwise distinct $f$-periodic points in $X(\overline{K})$. Assume that $(x_{n})_{n=1}^{\infty}$ is generic in $X$, i.e., every proper Zariski closed subset $Z\subsetneqq X$ contains only finitely many $x_{n}$. Then for every proper subvariety $V\subsetneqq X$, we have
(3.11) $\frac{\#(V\cap O_{f}(x_{n}))}{\#O_{f}(x_{n})}\to 0,\text{ as }n\to+\infty,$
@@
$\frac{\#(V\cap O_{f}(x_{n_{k_{l}}}))}{\#O_{f}(x_{n_{k_{l}}})}\to 0,\text{ as }l\to+\infty.$
[…]
@@
(3.12) $\int 1_{V}dm=0.$
@@
$\frac{\#(V\cap O_{f}(x_{n}))}{\#O_{f}(x_{n})}=\int 1_{V}dm_{n}\to\int 1_{V}dm,\text{ as }n\to+\infty.$
Suppose that (3.12) fails. Then there must be a point $y\in S$ (in the scheme-theoretic sense) with $a_{y}>0$ and $V\cap O_{f}(y)\neq\emptyset$. Let $k$ be the exact period of $y$ under $f$, and let $Y$ be the Zariski closure of $\{y\}$. Then $Y\subseteq\cup_{j=0}^{k-1}f^{j}(V)$, so $Y$ is a proper Zariski closed subset of $X$. Note that
$\frac{\#(Y\cap O_{f}(x_{n}))}{\#O_{f}(x_{n})}=\int 1_{Y}dm_{n}\to\int 1_{Y}dm\geq\frac{a_{y}}{k}>0,\text{ as }n\to+\infty.$
Hence for all sufficiently large $n\gg 1$, we have
@@
where the last equality follows from $f^{k}(y)=y$ and $Y=\overline{\{y\}}^{{\rm Zar}}$. Since
@@
the set $\cup_{j=0}^{k-1}f^{j}(Y)$ contains only finitely many $x_{n}$ because $(x_{n})$ is generic, contradicting (3.13). ∎
@@
Recall that $f:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ is a rational map of degree $d\geq 2$, and there is a number field $K$ such that for every $z_{0}\in{\rm Per}(f)$, we have $\lvert\rho_{f}(z_{0})\rvert^{n}\in K$ for some integer $n=n(z_{0})\geq 1$. We need to show that such an $f$ must be exceptional (Definition 1.1).
Assume by contradiction that $f$ is not exceptional. In particular, $f$ is not a flexible Lattès map. For every $z_{0}\in{\rm Per}(f)$, we have $\lvert\rho_{f}(z_{0})\rvert\in\overline{\mathbb{Q}}$, so $f$ is defined over $\overline{\mathbb{Q}}$ (hence over some number field) by Theorem 1.2. After replacing $K$ by a suitable finite extension, we may assume that both $f$ and $\overline{f}$ are defined over $K$, where $\overline{f}$ denotes the rational map obtained by conjugating the coefficients of $f$. According to [10, Theorem 7 and Lemma 9], which are essentially due to Zdunik [35], there exists a sequence $(x_{n})_{n=1}^{\infty}$ of distinct points in ${\rm Per}^{*}(f)$ such that
@@
Clearly, $\mathcal{L}_{f}=\mathcal{L}_{\overline{f}}$. For any $x\in{\rm Per}(f)$, we have
$\overline{x}\in{\rm Per}(\overline{f}),\quad n_{f}(x)=n_{\overline{f}}(\overline{x}),\quad\text{and}\quad\rho_{f}(x)=\overline{\rho_{\overline{f}}(\overline{x})},$
so $\chi_{f}(x)=\chi_{\overline{f}}(\overline{x})$.